From an 8-ounce (oz) bottle of medicine, how many doses can be obtained if each dose is two teaspoons (tsp)? (1 Tablespoon = 3 tsp = 0.5 oz).

16
24
36
48
Answer #1

Given, 8 ounce(oz) bottle of medicine.
1 oz = 29.574 ml
8 oz = 29.574×8 ml
=236.59 ml
given, each dose = 2 teaspoons
Now, 3 teaspoons= 0.5 oz
1 oz = 29.574 ml
0.5 oz=29.574×0.5 ml
=14.787 ml
So, 3 teaspoons = 14.787 ml
Now, using unitary method:-

14.787 ml = 3 teaspoons
1 ml = 3/14.787 teaspoons
236.59 ml = (3/14.787)×236.59 teaspoons
= 48 teaspoons
Now, 2 teaspoons = 1 dose
1 teaspoon= 1/2 dose
48 teaspoons= (1/2)×48 dose
=24 dose
Therefore, 24 dose can be obtained if each dose is two teaspoons.